lightfoot Inc., a software development firm, has stock outstanding as follows: 10,000 shares of cumulative preferred 3% stock, $20 par, and 13,000 shares of $50 par common. During its first four years of operations, the following amounts were distributed as dividends: first year, $2,300; second year, $3,800; third year, $22,430; fourth year, $41,750.

Calculate the dividends per share on each class of stock for each of the four years. Round all answers to two decimal places. If no dividends are paid in a given year, enter "0".

1st Year 2nd Year 3rd Year 4th Year
Preferred stock (dividend per share) $ $ $ $
Common stock (dividend per share)
Answer #1

Divided payable to preferred shareholders = 10,000 X $20 X 3% = $6,000

Dividend payable for 2 years to preferred shareholders = $6,000 X 2 = $12,000

Actual dividend paid to preferred shareholders = $2,300 + $3,800 = $6,100

Dividend payable to preferred shareholders third year = ($12,000 - $6,100) + $6,000 = $11,900

Dividend paid to common shareholders thirds year = $22,430 - $11,900 = $10,530

Dividend payable to preferred shareholders fourth year = $6,000

Dividend paid to common shareholders thirds year = $41,750 - $6,000 = $35,750

Dividend paid to Preferred shareholders per share:

1st year = $2,300 / 10,000 = $0.23

2nd year = $3,800 / 10,000 = $0.38

3rd year = $11,900 / 10,000 = $1.19

4th year = $6,000 / 10,000 = $0.60

Dividend paid to common shareholders per share:

1st year = $0

2nd year = $0

3rd year = $10,530 / 13,000 = $0.81

4th year = $35,750 / 13,000 = $2.75

1st Year 2nd Year 3rd Year 4th Year
Preferred stock (dividend per share) $0.23 $0.38 $1.19 $0.60
Common stock (dividend per share) $0 $0 $0.81 $2.75